[Brahms-l] Shift report 20050309 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Mar 09 2005 - 08:05:02 EST
Shifter leader was Sorin ZGURA
Shifters were 


1. Machine status to the entry of shift:
00:00 RHIC INJECTION POWER SUPPLY PROBLEM!
00:50-  EXPERIMENT POWER SUPPLY PROBLEM 
           STAR magnet power supply problem 
05:40: Working on ramp for low-energy Cu run.

2. No runs. 

3. No detector problems! 


 

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Mar 9 08:05:30 2005

This archive was generated by hypermail 2.1.8 : Wed Mar 09 2005 - 08:05:45 EST